# Universal Approximation Theory: The Basic Theory for Transformer-based Large Language Models

### v3

@@ Abstract
Language models have become a focal point in the field of artificial intelligence, especially following the emergence of ChatGPT. Large-scale Transformer networks have quickly become the primary method for advancing natural language processing algorithms. These models, built on the Transformer architecture, are capable of engaging in near-human-like interactions and can even assist in guiding human tasks. Despite their impressive capabilities, there remain some critical theoretical questions in the development of large language models (LLMs): What makes the Transformer architecture so effective in driving intelligent language applications, such as translation and coding? What underlies the in-context learning (ICL) capability of LLMs? How does the LoRA approach enhance the fine-tuning of LLMs? What supports the feasibility of pruning in LLMs? To address these key questions and explore the technical strategies within LLMs, we draw on the Universal Approximation Theory (UAT) to provide a theoretical foundation, revealing the mechanisms behind these advancements.
@@ 1 Introduction
In recent years, the rapid emergence of LLMs in the field of artificial intelligence has undoubtedly become one of the most notable advancements within the domain. The core allure of these models stems from their extraordinary capabilities in language processing. Language, as a unique crystallization of human intelligence, serves not only as the external reflection of thought but also as the bridge for communication, the cornerstone for the dissemination of knowledge, and the continuation of civilization, profoundly shaping the identity of humans as a unique species. Thus, endowing machines with the ability to understand and generate language marks a significant leap towards the realization of true artificial intelligence. The emergence of models such as the ChatGPT (Radford and Narasimhan 2018; Brown et al. 2020; Achiam et al. 2023), the Llama (Touvron et al. 2023), and the PaLM (Chowdhery et al. 2023) vividly demonstrates this point.
@@
A distinctive feature of LLMs is their immense parameter size (Achiam et al. 2023; Touvron et al. 2023; Chowdhery et al. 2023; Chen et al. 2021; Zeng et al. 2022), often amounting to hundreds of billions or even trillions (for instance, GPT-3’s (Brown et al. 2020) 175 billion parameters and PaLM’s (Chowdhery et al. 2023) 540 billion parameters). This vast parameter scale lays the foundation for their exceptional language processing capabilities and enables them to exhibit almost human-like traits, such as ICL  (Brown et al. 2020; Dong et al. 2022), instruction following (Sanh et al. 2021; Ouyang et al. 2022; Wei et al. 2021), and multi-step reasoning (Wei et al. 2022). Notably, these colossal models are predominantly trained by tech giants like Google and Microsoft using large-scale GPU clusters (Zhao et al. 2023), sparking a research fervor on how to efficiently fine-tune them with limited GPU resources. The advent of Lora (Hu et al. 2021) fine-tuning technology has provided an effective pathway for this, allowing for the fine-tuning of large models under resource constraints without the need to adjust all parameters of the original model comprehensively. Moreover, model pruning techniques (Sun et al. 2023; Ma, Fang, and Wang 2023) are crucial for deploying large models in resource-constrained environments, aiming to reduce the model size for operation on smaller devices. Today, LLMs possess a diverse range of functionalities, from translation and text summarization to automatic code generation, demonstrating their versatility.
@@
Despite the rapid advancements of LLMs towards higher intelligence and reliability, the theoretical foundations behind these models remain largely unexplored and shrouded in uncertainty. The scientific community is actively seeking to uncover the intrinsic mechanisms behind their powerful capabilities, including the analysis of ICL mechanisms (Xie et al. 2021; Min et al. 2022). In this context, we establish the UAT (Cybenko 2007; Hornik, Stinchcombe, and White 1989) as the mathematical essence of Transformer-based LLMs and use the perspective of UAT to elucidate key technologies and phenomena within LLMs, aiming to provide explanations from a new angle.
@@
While previous studies have explored the relationship between Transformers and UAT (Yun et al. 2019, 2020; Kratsios et al. 2021; Alberti et al. 2023), these studies are based on a series of complex assumptions and derivations, lacking generality and failing to intuitively address some of the issues present in LLMs. In contrast, this paper aims to provide proof that bridges Transformers and UAT, grounded in the foundational mathematical form of UAT as established by Cybenko; Hornik, Stinchcombe, and White. Our proof is characterized by its simplicity, absence of assumptions, and generality, making it applicable to Convolutional Neural Networks (CNNs) as well (Wang and Li 2024b). Our approach not only provides intuitive answers to some of the challenges within LLMs but also aids in the design of deep learning networks from a theoretical standpoint (Wang and Li 2024a). These capabilities are not addressed by previous theories. Our contributions are as follows:
  • We prove that the Transformer is the tangible embodiment of UAT.
@@
  • We deliver a rigorous scientific explanation of Transformer-based LLMs through the UAT.
@@
Our article is structured as follows: In Section 2, we begin by introducing the UAT and propose that to demonstrate the Transformer’s adherence to UAT, it is necessary to show that both Linear and Multi-Head Attention (MHA) can be represented in the form of matrix-vector multiplication. In Section 3, we establish that the Transformer belongs to the category of UAT. Our proof strategy unfolds starting with Section 3.1, where we present the idea to express Linear and MHA as matrix-vector products. In Section 3.2, we detail the computational methods required for the proof. Sections 3.3 and 3.4 are dedicated to proving that both Linear and MHA can indeed be represented as matrix-vector operations. Further in Section 4, we leverage UAT to theoretically elucidate some fundamental issues (Generalization 4.1, ICL 4.2) and techniques (Pruning 4.3 and LoRA 4.4) associated with LLMs. Finally, we conclude with a summary of the existing challenges and potential future developments for LLMs in Section 4.5.
2 The Universal Approximation Theory
Up to now, the UAT (Cybenko 2007) is the most widely recognized fundamental theory in deep learning. However, this theorem is only applicable to the simplest form of neural networks, the multilayer perceptron (MLP) (Cybenko 2007; Popescu et al. 2009). Due to the increased complexity of Transformer networks, they cannot be mathematically expressed in the same form as UAT. Thus, the theory has not yet been extended to Transformer networks. Our aim in this paper is to unify Transformer networks under the framework of UAT, thereby standardising their mathematical representation. Before unifying their mathematical forms, we provide a brief overview of the theorem, which was initially proposed by Cybenko (2007). This theorem encompasses numerous conclusions and proof details. Although it has been further developed, its fundamental mathematical form remains unchanged. Therefore, this paper explains the theory based on the UAT form presented by Cybenko (2007). Theorem 2 from Cybenko (2007) states that if $\sigma$ is any continuous sigmoidal function, then finite sums of the following form:
$G(\mathbf{x})=\sum_{j=1}^{N}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ (1)
@@ 2 The Universal Approximation Theory
is dense in $C\left(\mathbf{I}_{n}\right)$. Here, $\mathbf{W}_{j}\in\mathbb{R}^{n}$ and $\alpha_{j},\theta\in\mathbb{R}$ are fixed. For any $f\in C\left(\mathbf{I}_{n}\right)$ and $\varepsilon>0$, there exists a function $G(\mathbf{x})$:
$|G(\mathbf{x})-f(\mathbf{x})|<\varepsilon\quad\text{ for all }\quad\mathbf{x}\in\mathbf{I}_{n}.$ (2)
@@
This implies that, when $N$ is sufficiently large, a neural network can approximate any continuous function on a closed interval. Hornik, Stinchcombe, and White (1989) further demonstrates that multilayer feedforward networks also conform to the UAT, capable of approximating arbitrary Borel measurable functions. Observing Equation (1), where the function $G(\mathbf{x})$ yields a scalar output in $\mathbb{R}$, the scenario expands naturally when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring the approximation in each dimension. It becomes evident that to accommodate this multidimensional output, a simple extension to Equation (1) suffices: the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix. Nevertheless, in these formulas, the theorem does not straightforwardly apply to Transformer architectures. However, if we succeed in reformulating both the Linear ($\mathbf{x}|\mathbf{W}$) and MHA ($\mathbf{x}|\mathbf{W}$) components into a unified representation, $\mathbf{W}^{\prime}\mathbf{x}^{\prime}$, where $\mathbf{W}$ and $\mathbf{x}$ represent the parameters and inputs for each component respectively, and $\mathbf{W}^{\prime}$ is derived from $\mathbf{W}$ while $\mathbf{x}^{\prime}$ signifies a column vector derived from $\mathbf{x}$ (for details, refer to Section 3.1). It is easy to prove that multi-layer Transformer networks are concrete embodiments of UAT.
Consequently, our objective is to represent Linear and MHA architectures as a matrix-vector multiplication similar to Eq. (3). This approach offers two key benefits: Firstly, it allows for a fundamental mathematical investigation of deep learning models. Secondly, it provides a more intuitive and objective basis for comparing differences among various models. On this foundation, we propose to transform Linear and MHA into the matrix-vector format.
3 UAT for Transformer
In the previous section, we have outlined our fundamental goal to unify the Transformer model under the framework of UAT by converting Linear and MHA operations in the Transformer into matrix-vector forms. In this section, we will use the Matrix-Vector method to transform Transformer operations into matrix times vector forms.
3.1 Matrix-Vector Method
@@
Before delving into transforming Linear and MHA into their matrix-vector format, we introduce the Matrix-Vector Method, which will subsequently be employed to cast both Linear and MHA operations into a unified matrix-vector format. This method constitutes a strategic realignment of input data and corresponding parameters of various transformations within the network, as illustrated in Figure 1. The underlying principle is as follows: both the input ($\mathbf{x}$) and output ($\mathbf{y}$) data are reconfigured uniformly through a transformation $\mathbf{T_{D}}$ into column vectors ($\mathbf{x}^{\prime}$ and $\mathbf{y}^{\prime}$, while parameter tensor ($\mathbf{W}$) is reorganized into matrix form ($\mathbf{W}^{\prime}$). A critical requirement is that identical operations throughout the LLMs network adhere to the same restructuring scheme, thereby eliminating the need for additional transformations on intermediate feature data; they can directly be represented as column vectors. Additionally, by default, matrix variables in the original formulas are represented in bold, such as $\mathbf{x}$, and in the Matrix-Vector form, corresponding variables are denoted with a prime symbol () in the upper right corner, such as $\mathbf{x}^{\prime}$. Elements within matrices are represented by corresponding lowercase letters with subscripts, for example, $x_{i}$. The Matrix-Vector Method can be succinctly encapsulated as follows:
$\mathbf{y}=T(\mathbf{x}|\mathbf{W})\rightarrow\mathbf{y}^{\prime}=\mathbf{W}^{\prime}\mathbf{x}^{\prime}$ (3)
@@
It is obvious that $T_{D}$ and $T_{P}$ are not fixed, we could design various kinds of ways to do those. For convenience, we present a methodology tailored for Linear in Section 3.3 and an approach for MHA in Section 3.4, thereby illustrating the adaptability and application of the Matrix-Vector Method across different components of the Transformer architecture.
3.2 Diamond Matrix Multiplication
While our goal is to transform Linear and MHA modules into a matrix-vector format, this process is complex. In order to describe this transformation process clearly, we propose a new computational method called the Diamond Multiplication Method, denoted by the symbol $\diamond$. (The purpose of proposing this method is to clearly illustrate the calculation process in turning MHA into Matrix-Vector format. If we use direct matrix multiplication, some computational details cannot be visually distinguished.)
$\left(\begin{array}[]{cccc}w_{1,1}&w_{1,2}&\cdots&w_{1,n}\\ w_{2,1}&w_{2,2}&&w_{2,n}\\ \vdots&\vdots&\vdots&\vdots\\ w_{m,1}&w_{m,2}&&w_{m,n}\end{array}\right)\diamond\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ \vdots\\ x_{m}\end{array}\right)$ (4)
$=\left(\begin{array}[]{c}w_{1}x_{1}+w_{21}x_{2}+\cdots w_{m,1}x_{m}\\ w_{1,2}x_{1}+w_{2,2}x_{2}+\cdots w_{m,2}x_{m}\\ \vdots\\ w_{1,n}x_{1}+w_{2,n}x_{2}+\cdots w_{mn}x_{m}\end{array}\right)$
Let $\mathbf{W}\in\mathbb{R}^{(m,n)}$ and $\mathbf{x}\in\mathbb{R}^{(m,1)}$, the Diamond Matrix Multiplication is defined as $\mathbf{y}=\mathbf{W}\diamond\mathbf{x}$, where $\mathbf{y}\in\mathbb{R}^{(m,1)}$. The computation procedure involves element-wise multiplication and summation of corresponding elements from left to right across the columns of matrix $\mathbf{W}$ with vector $\mathbf{x}$. The sum of the multiplication of the $i$th column of $\mathbf{W}$ with $\mathbf{x}$ is assigned as the $i$th element of $\mathbf{y}$. A detailed calculation process is provided in Eq. (4). The computation process for $\mathbf{y}=\mathbf{x}\diamond\mathbf{W}$ is identical to that of $\mathbf{W}\diamond\mathbf{x}$, implying that $\mathbf{W}\diamond\mathbf{x}=\mathbf{x}\diamond\mathbf{W}$.
Furthermore, based on its computation method, it can be inferred that the Diamond Matrix Multiplication is related to conventional matrix multiplication, specifically: $\mathbf{W}\diamond\mathbf{x}=\mathbf{W}^{T}\mathbf{x}$. Additionally, when $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are square matrices, we can derive the following relationships: $\mathbf{W}_{1}\diamond[\mathbf{x}\diamond\mathbf{W}_{2}]=\mathbf{W}_{2}^{T}\diamond\mathbf{W}_{1}\diamond\mathbf{x}$ and $[\mathbf{W}_{1}\diamond\mathbf{x}]\diamond\mathbf{W}_{2}=\mathbf{W}_{1}^{T}\diamond\mathbf{W}_{2}\diamond\mathbf{x}$. Detailed derivations can be found in Appendix A.

### v4

# Dynamic Universal Approximation Theory: The Basic Theory for Transformer-based Large Language Models

@@ Abstract
Language models have become a focal point in the field of artificial intelligence, especially following the emergence of ChatGPT. Large-scale Transformer networks have quickly become the primary method for advancing natural language processing algorithms. These models, built on the Transformer architecture, are capable of engaging in near-human-like interactions and can even assist in guiding human tasks. Despite their impressive capabilities, there remain some critical theoretical questions in the development of large language models (LLMs): What makes the Transformer architecture so effective in driving intelligent language applications, such as translation and coding? What underlies the in-context learning (ICL) capability of LLMs? How does the LoRA approach enhance the fine-tuning of LLMs? What supports the feasibility of pruning in LLMs? To address these key questions and explore the technical strategies within LLMs, we draw on the Universal Approximation Theory (UAT) and then propose the Dynamic Universal Approximation Theory (DUAT) to provide a theoretical foundation, revealing the mechanisms behind these advancements.
@@ 1 Introduction
In recent years, the rapid emergence of LLMs in the field of artificial intelligence has undoubtedly become one of the most notable advancements within the domain. The core allure of these models stems from their extraordinary capabilities in language processing. Language, as a unique crystallization of human intelligence, serves not only as the external reflection of thought but also as the bridge for communication, the cornerstone for the dissemination of knowledge, and the continuation of civilization, profoundly shaping the identity of humans as a unique species. Thus, endowing machines with the ability to understand and generate language marks a significant leap towards the realization of true artificial intelligence. The emergence of models such as the ChatGPT Radford & Narasimhan (2018); Brown et al. (2020); Achiam et al. (2023), the Llama Touvron et al. (2023), and the PaLM Chowdhery et al. (2023) vividly demonstrates this point.
@@
A distinctive feature of LLMs is their immense parameter size Achiam et al. (2023); Touvron et al. (2023); Chowdhery et al. (2023); Chen et al. (2021); Zeng et al. (2022), often amounting to hundreds of billions or even trillions (for instance, GPT-3’s Brown et al. (2020) 175 billion parameters and PaLM’s Chowdhery et al. (2023) 540 billion parameters). This vast parameter scale lays the foundation for their exceptional language processing capabilities and enables them to exhibit almost human-like traits, such as ICL  Brown et al. (2020); Dong et al. (2022), instruction following Sanh et al. (2021); Ouyang et al. (2022); Wei et al. (2021), and multi-step reasoning Wei et al. (2022). Notably, these colossal models are predominantly trained by tech giants like Google and Microsoft using large-scale GPU clusters Zhao et al. (2023), sparking a research fervor on how to efficiently fine-tune them with limited GPU resources. The advent of Lora Hu et al. (2021) fine-tuning technology has provided an effective pathway for this, allowing for the fine-tuning of large models under resource constraints without the need to adjust all parameters of the original model comprehensively. Moreover, model pruning techniques Sun et al. (2023); Ma et al. (2023) are crucial for deploying large models in resource-constrained environments, aiming to reduce the model size for operation on smaller devices. Today, LLMs possess a diverse range of functionalities, from translation and text summarization to automatic code generation, demonstrating their versatility.
@@
Despite the rapid advancements of LLMs towards higher intelligence and reliability, the theoretical foundations behind these models remain largely unexplored and shrouded in uncertainty. The scientific community is actively seeking to uncover the intrinsic mechanisms behind their powerful capabilities, including the analysis of ICL mechanisms Xie et al. (2021); Min et al. (2022). In this context, we leverage UAT and then establish the DUAT as the mathematical essence of Transformer-based LLMs and use the perspective of DUAT to elucidate key technologies and phenomena within LLMs, aiming to provide explanations from a new angle.
@@
Although previous studies have explored the relationship between Transformer Vaswani2017AttentionIA and the UAT Yun et al. (2019; 2020); Kratsios et al. (2021); Alberti et al. (2023), they rely on a series of complex assumptions and derivations, which limits their generalizability and fails to address certain issues in LLMs in a straightforward manner. In contrast, this paper directly demonstrates the connection between multi-layer Transformer and UAT by representing it in a mathematical format. Cybenko (1989) and Hornik et al. Hornik et al. (1989) established the fundamental mathematical form of UAT, which we extend to the DUAT by leveraging the residual structure of Transformers. This form can account for several key issues in LLMs, such as why LLMs are capable of handling both translation and generation tasks simultaneously, and whether LLMs possess memory capacity (Wang & Li, 2024a). Moreover, DUAT is equally applicable to convolutional neural networks (CNNs) Wang & Li (2024b). Our approach not only provides intuitive explanations for some of the challenges within LLMs but also offers theoretical insights that aid in the design of deep learning networks Wang & Li (2024c), covering capabilities that previous theories have not addressed. Our contributions are as follows:
  • We further develop the UAT to DUAT.
  • We prove that the Transformer is the tangible embodiment of DUAT.
@@
  • We deliver a rigorous scientific explanation of Transformer-based LLMs through the DUAT.
@@
The structure of this paper is as follows: in Section 2.1, we first introduce the UAT and propose that to prove Transformers adhere to the DUAT framework, it is necessary to demonstrate that both the linear layer and the multi-head attention mechanism (MHA) can be expressed in the form of matrix-vector multiplication. Then, in Section 2.2, we present a general framework for representing various operations in deep learning as matrix-vector multiplication. In Section 3, we establish that Transformers fall within the scope of DUAT. Our proof strategy involves representing the linear layer and MHA as matrix-vector operations. Sections 3.1 and 3.2 provide detailed proofs that the linear layer and MHA can indeed be expressed in this form. In Section 3.3, we present the DUAT representation of a multi-layer Transformer network. In Section 4, we leverage the DUAT framework to theoretically elucidate several fundamental questions (e.g., generalization ability, 4.1, and in-context learning capability, 4.2) and technical aspects (e.g., feasibility of pruning, 4.3, and the effectiveness of the LoRA scheme, 4.4) related to large language models (LLMs). Finally, in Section 6, we summarize the current challenges of LLMs and explore potential directions for future research.
2 The Universal Approximation Theory and Matrix-Vector Method
2.1 The Universal Approximation Theory
The Universal Approximation Theorem (UAT) (Cybenko, 1989) remains one of the most widely recognized foundational theories in the field of deep learning to date. However, the theorem primarily applies to the simplest forms of neural networks—single-layer or multilayer perceptrons (MLPs) (Cybenko, 1989; Popescu et al., 2009). Due to the significantly increased complexity of Transformer networks, they cannot be expressed mathematically in the same manner as those governed by UAT. Consequently, the extension of UAT to Transformer networks has yet to be established.
This paper aims to develop the DUAT (an advance of UAT) and explore Transformer networks within this framework to standardize their mathematical representation. Before achieving this, we will briefly revisit the UAT as initially proposed by Cybenko (1989). The theorem encompasses a wealth of conclusions and detailed proofs. Although it has been further developed over time, its core mathematical form remains unchanged. Thus, this paper interprets the theory based on the UAT form articulated by Cybenko (1989).
According to Theorem 2 in Cybenko (1989), if $\sigma$ is any continuous sigmoid function, then a finite sum of the following form:
$G(\mathbf{x})=\sum_{j=1}^{N}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+b_{j}\right)$ (1)
@@ 2.1 The Universal Approximation Theory
is dense in $C\left(\mathbf{I}_{n}\right)$. Here,$\mathbf{x}\in\mathbf{I}_{n}$, $\mathbf{W}_{j}\in\mathbb{R}^{n}$ and $\alpha_{j},b\in\mathbb{R}$ are fixed. For any $f\in C\left(\mathbf{I}_{n}\right)$ and $\varepsilon>0$, there exists a function $G(\mathbf{x})$:
$|G(\mathbf{x})-f(\mathbf{x})|<\varepsilon$ (2)
@@
This implies that, when $N$ is sufficiently large, a neural network can approximate any continuous function on a closed interval. Hornik et al. (1989) further demonstrates that multilayer feedforward networks also conform to the UAT, capable of approximating arbitrary Borel measurable functions. Observing Eq. 1, where the function $G(\mathbf{x})$ yields a scalar output in $\mathbb{R}$, the scenario expands naturally when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring the approximation in each dimension. It becomes evident that to accommodate this multidimensional output, a simple extension to Eq. 1 suffices: the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $b_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix. Nevertheless, in these formulas, the theorem does not straightforwardly apply to Transformer architectures. However, if we succeed in reformulating both the Linear ($\mathbf{x}|\mathbf{W}$) and MHA ($\mathbf{x}|\mathbf{W}$) components into a unified representation, $\mathbf{W}^{\prime}\mathbf{x}^{\prime}$, where $\mathbf{W}$ and $\mathbf{x}$ represent the parameters and inputs for each component respectively, and $\mathbf{W}^{\prime}$ is derived from $\mathbf{W}$ while $\mathbf{x}^{\prime}$ signifies a column vector derived from $\mathbf{x}$. It is easy to build the connection between multi-layer Transformer networks and UAT.
2.2 The Matrix-Vector Method
@@
Before delving into transforming Linear and MHA into their matrix-vector format, we introduce the Matrix-Vector Method, which will subsequently be employed to cast both Linear and MHA operations into a unified matrix-vector format. This method constitutes a strategic realignment of input data and corresponding parameters of various transformations within the network, as illustrated in Figure 1. The underlying principle is as follows: both the input ($\mathbf{x}$) and output ($\mathbf{y}$) data are reconfigured uniformly through a transformation $\mathbf{T_{D}}$ into column vectors ($\mathbf{x}^{\prime}$ and $\mathbf{y}^{\prime}$, while parameter tensor $\mathbf{W}$ is reorganized into matrix form $\mathbf{W}^{\prime}$ by $\mathbf{T_{P}}$. $\mathbf{T_{D}}$ and $\mathbf{T_{P}}$ are transformations and their purpose is to reorganize the input $\mathbf{x}^{\prime}$ and $\mathbf{y}^{\prime}$ into column vectors and $\mathbf{W}^{\prime}$ fulfill $\mathbf{W}^{\prime}\mathbf{x}^{\prime}=\mathbf{y}^{\prime}$. A critical requirement is that the input and output throughout the LLMs network adhere to the same restructuring scheme, thereby eliminating the need for additional transformations on intermediate feature data; they can directly be represented as column vectors. Additionally, by default, matrix variables in the original formulas are represented in bold, such as $\mathbf{x}$, and in the Matrix-Vector form, corresponding variables are denoted with a prime symbol () in the upper right corner, such as $\mathbf{x}^{\prime}$. Elements within matrices are represented by corresponding lowercase letters with subscripts, for example, $x_{i}$. The Matrix-Vector Method can be succinctly encapsulated as follows:
$\mathbf{y}=T(\mathbf{x}|\mathbf{W})\mapsto\mathbf{y}^{\prime}=\mathbf{W}^{\prime}\mathbf{x}^{\prime}$ (3)
@@
where $T$ represents a general transformation, like MHA. It is obvious that $T_{D}$ and $T_{P}$ are not fixed, we could design various kinds of ways to do those. For convenience, we present a methodology tailored for Linear in Section 3.1 and an approach for MHA in Section 3.2, thereby illustrating the adaptability and application of the Matrix-Vector Method across different components of the Transformer architecture.
Figure 2: This diagram shows the differences between an UAT and a DUAT.
2.3 The introduction of DUAT
Although the UAT function has strong approximation capabilities, it has a fatal flaw: it can only fit one function within a closed interval. This limitation results in poor generalization.Therefore, to enable the model to fit different functions within a closed interval based on the input, we need to grant the parameters in the UAT the ability to dynamically adjust according to the input, as defined in Eq. 4.
$\mathbf{y}_{i}={G(\mathbf{x}_{i}|\mathbf{W(\mathbf{x}_{i})},\mathbf{b(\mathbf{x}_{i})})}$ (4)
where $\mathbf{W}=\mathbf{W}_{1}\cdots\mathbf{W}_{N}$ and $\mathbf{b}=b_{1}\cdots b_{N}$igure 2 illustrates the comparison between UAT and DUAT. On the left side, Figure 2.a demonstrates the deficiency in data recognition. The input to the network is shown, with only the initial part differing slightly. This can be likened to tasks in LLMs, such as translating, continuing, or summarizing the same passage, where only the initial text segment differs while the rest remains the same. Figure 2.b provides a schematic illustration of the output for the input shown in Figure 2.a. Because the UAT can only fit a single function, the output closely follows the trend of the input, resulting in nearly identical outcomes regardless of the slight differences in the initial part. Essentially, the output remains the same regardless of the prompt.
In contrast, Figure 2.c shows the schematic output of the DUAT (Dynamic Universal Approximation Theorem) function. In DUAT, some or all parameters are functions of the input, allowing it to produce different outputs based on different prompts. This flexibility enables DUAT to generate varied results.
On the right side, Figure 2 compares UAT and DUAT through the distribution of data. Since the parameters of UAT are fixed, it can only fit a fixed data distribution. Given that we typically train with large datasets, the goal for UAT is to fit as much data as possible, which often results in fitting only the densest part of the data distribution, represented in purple. On the other hand, DUAT can dynamically adjust its parameters based on the input, allowing it to fit multiple different distributions simultaneously.
This comparison highlights the advantages of DUAT in handling more complex and diverse data scenarios, making it a more versatile and powerful tool for various applications.
3 DUAT for Transformer
In this section, we will apply the matrix-vector approach to convert the operations in the Transformer Vaswani2017AttentionIA into matrix-vector multiplication form, and then explain the defination of DUAT and demonstrate that the Transformer is essentially a DUAT function.